Please see end of minutes.
Rob gave an overview of the project: The Chemistry Library closed at the end of August. All collections were to be removed and the space returned to the Chemistry Department, which would use it for office space needed for a building addition. Collections were to be transferred to either ANNEX or LSM in September.
While DTS was weeding and Access Services was shelf-reading, a mold outbreak was discovered in CHEM REF and PER. Affected items were pulled and work orders were placed to have Facilities look at the problem. However, RUL had to put the moving contract out to bid, which extended the project timeline.
In the interim, the malfunctioning HVAC system started a larger mold outbreak, which was not caught because the building was not in use. At an unknown point the HVAC system was fixed, which stabilized the environment and prevented the mold from spreading further. The project parameters changed to a major weeding/removal/remediation project. A call went out to Access Services for volunteers to help with the weeding/removal. Three 30 cubic yard dumpsters were filled with materials that were weeded from the collection. Remaining material was then packed by Library Interiors and is being remediated by Rapid Refile.
A discussion regarding mold and the appropriate response ensued. Anthony researched the issue and contacted colleagues at several cultural institutions to see how they dealt with it. All recommended a higher level of personal protection (best practices) than that recommended by Rutgers University's Environmental Health and Safety office (REHS). They also noted that large outbreaks should be handled by professional contractors. Andy noted that since REHS cleared the space for human occupation (the outbreak was less than 10 linear feet), then the higher protection standard was not necessary. The project was successfully completed by the deadline, the end of the first full week of November.
Andy gave us background on how RUL has been monitoring conditions (temperature and humidity) inside our buildings. It has been a combination of using digital data loggers, a hand-held thermometer/hygrometer, and personal observation. RUL's loggers are at least ten years old and the interface is obsolete. Many of our current data loggers no longer function. Melanie discussed how using a data logger prevented a problem from developing at the Math Library by proving information that allowed her to alert Facilities to a ventilation issue.
Andy and Rob noted that RUL have little direct control over the conditions inside their buildings: Heating, ventilation, and cooling (HVAC) are mostly controlled by Facilities. Therefore, any changes or problems must be reported via a work order. Also, Bill Puglisi, RUL's buildings coordinator, could assist as well. Andy recommended that Unit Coordinators or Supervisors should be responsible for maintaining future environmental records rather than have one person collect all of the data and act on it after the fact.
Andy and Rob passed around handouts on data logger and hand held laser thermo-hygrometer pricing and specifications from several vendors.
There was a discussion on what types of data could be output from newer models of data loggers. There was also a discussion of what interval the data should be collected at. The current HOBO data loggers record every sixty seconds, which is a large amount of information to deal with. The group found that the newer models allowed for customized data recordings. Andy noted that having laser thermo-hygrometers would allow Collection Management personnel to spot check for potential problems in the collections prior to obtaining report data and would be useful for preventing problems quickly.
Melanie recommended looking at the Image Permanence Institute, part of the Rochester Institute of Technology, which offers services for data analysis of temperature and humidity readings. The IPI offers several levels of service, which is priced based on the size of the institution requesting analysis.
Jimmie mentioned that Smith Library has no problems regarding mold. He also noted that the Physical Plant Center adjusts the environmental conditions and that there is nothing for the staff to control.
Rob discussed the data loggers at ALEX and the building's micro-environments. Bill asked where/how they are attached? They are velcroed to the top of the stacks-out of sight of patrons. Discussion ensued regarding how to come up with averages for temperature and humidity.
Andy moved that we form a sub-group to look at new data loggers as well as a hand-held temperature/hygrometer device. The sub-group will put together a requirements list which will determine the best products. The group includes Andy, Rob, Melanie, and Anthony. Meeting time TBD.
Rob gave an overview of the space management database, created in Microsoft Excel. The database will allow Collection Management to better distribute remaining free space in a collection, provide a tool to help selectors identify space issues and allow for the tracking of transfers between libraries in the event of weeding or consolidation projects. The new space management database will replace the shift measurement form that was previously used to calculate large collection shifts.
Rob explained that the new space management database had several advantages over the old shift calculations. The new database takes into account materials that are currently checked out by patrons, automatically adjusts the remaining space as materials are added to the collection and would eliminate the need to measure a collection prior to initiating a shifting project. To do this, the database would be recalibrated every time a major shifting project occurred.
All CMG members were e-mailed a copy of the database and a draft of instructions. Rob proceeded to explain the process and results using his test run on ALEX 2A. There will be several training sessions in the future with the goal of setting up space management databases in each library's collections.
Discussion among the group began with how to handle oversize items. Rob informed the group that it is up to CM staff to make the decision of what is an oversize item using established guidelines. Jimmie noted that binders will be problematic.
LSM and Math were both nominated as candidates for the implementing the next space management databases. LSM (Anthony) will have to coordinate with the DTS (Jamie) weeding project. The evening student assistants can assist with the measuring.
Irina observed that CM staff may be needed at libraries without permanent full -time staff to set up the databases.
If DTS pulls items in the future, then they will have to update the database for that particular library so that entries can be made into the transfer section of the database. Rob noted that only the number of materials being transferred, not the linear inches of the materials, needed to be reported; the database will extrapolate the number of inches based on statistical sampling. Jamie will need to inform the appropriate DTS staff so that this information can be updated.
A webinar called "Co-operative Community Emergency Response and Recovery Town Hall" is to be held Friday, December 6. CMG members interested were provided a web link to sign up. It is free and those who attend will provide an overview to the group.
The suspension of the bindery and the effects on current periodicals will be discussed.
Co-chair Robert Krack ended the meeting at 4:05 P.M.
ALEX: At Alexander we have begun expanding the Adopt-a-Shelf program to include not only students trained in Collection Management, but all student workers who have an available hour to work on the project. Both the feedback from the students in the program and the results have been positive so far. Many of the Collection Management trained students have requested a second area to maintain in addition to their original range. Work on the periodical project to incorporate transfers from Douglass has resumed now that work on the space management prototype has been completed. There are less than 2 sections of books remaining to be inserted and more space is being created for growing titles through the addition of more shelves. We are also performing some mini-shifts on the third floor and putting a few hours into the vacuuming project.
DOUGLASS: Here at Douglass our main focus continues to be the basement level shift of the A-L collection. Things are progressing slowly with the newer student workers as we are continuously shelf-reading the shifted areas to ensure no major mistakes occur during the process. We've moved about 3 ranges so far, and they have been labeled with call number guides. Earlier this week we also began shelving the new V/DISC (laser disc) collection into the tapestry room. These items will be shelved between the PN STACKS items and the bound PER items in that room. Each time ~2 shelves worth of discs are cataloged in-house, we will be moving them to the appropriate area and shelving them.
KILMER: Kilmer has no new updates since we last met. Most of our students are call number trained and up to date with the library policies. Shelving returns has been fine so far and shelf reading has started again for each month. The only negative news has been the increase in theft lately at Kilmer, mostly of Apple electronics.
LSM: LSM began lending laptops to patrons, which seems to be going well: none have been reported missing or damaged thus far. LSM is continuing the shelf-reading project of the second floor stacks. Returns are relatively tame. Lots of mini-shifting and one medium shift as certain parts of the collection are quite full. DTS began its weeding project in the second floor stacks. Ref weeding is winding down. Carpet tiles were installed on the first floor in two phases: first the public areas and then the conference room (matching those on the third floor). DTS:
LSM REF: Besides some minor follow up, LSM REF weeding should be complete at this point.
LSM STACKS: Selector decisions for LSM STACKS H1 - HA4585 call range have been completed. Lists were sent to the other library units for validations. All validations should be completed and the report ready for pulling at LSM by the week of 11/18th-22nd. New reports for LSM STACKS A, C & D call ranges have already been run and distributed to LSM selectors for decisions. Also the Chester collection is being transferred from LSM's second floor to the Annex. There are about 324 titles, of which only 17 are in the catalog. DTS staff are creating brief records for items not found in catalog with an NBINV note put in the bib record before transferring.
CHEM MOLD: Chemistry withdraw lists were distributed. All items have been pulled for withdraw and all remaining items for remediation have been packed by Libraries Interiors, now awaiting pick-up by Rapid Refile (the company that will remediate the items)
OTHER: A few miscellaneous print titles remain from the Douglass Periodicals consolidation project, DTS will resume work on those, as well as lots of microform titles that still need to be processed. It appears binding may be suspended indefinitely, which also means a discussion will be required here about how to house loose periodicals in the collection. More information should be available soon, topic should be added to future agenda (Dec./Jan. meeting?) And of course all routine work continues.